If you have 56 grams of O2 that react with NH3, how many moles H20 will be produced?


1
Expert's answer
2021-05-03T07:58:36-0400

4NH3 + 3O2 = 6H2O + 2N2

n = m / M

M (O2) = 32 g/mol

M (H2O) = 18 g/mol

n (O2) = 56 / 32 = 1.75 mol

According to the equation, n (H2O) = 6/3 x n(O2) = 2 x 1.75 = 3.5 mol